For its most recent fiscal year, Carmichael Hobby Shop recorded EBITDA of $512,725.20, EBIT of $362,450.20, zero interest expense, and cash flow to investors from operating activity of $348,461.25. Assuming there are no non-cash revenues recorded on the income statement, what is the firm's net income after taxes?
Answer #1

rate positively ..

Given that -
i EBITDA 512,725.20
ii EBIT 362,450.20
iii=i-ii Depreciation = 150,275.00
iv Cash flow to investors from operating activity 348,461.25
v=iv+iii Net income = 498,736.25
therefore, Answer = $ 498,736.25